OpenStudy (anonymous):

-7/8x=-5 how do you solve it can someone show me so I can do the rest on my on. thx

jhonyy9 (jhonyy9):

multiply both sides by 8 and will get -7x=-40 divide both sides by -7 and will get x= -40/(-7) = 40/7 x=40/7